OpenStudy (anonymous):

7th grade math help please

OpenStudy (anonymous):

@aha.its.kristen

OpenStudy (anonymous):

@Secret-Ninja

OpenStudy (anonymous):

@walter20

OpenStudy (anonymous):

@themaster02

OpenStudy (anonymous):

Whats the question?

OpenStudy (anonymous):

Carly will enter two baking competitions. The first competition requires her to make 2x + 4 cupcakes. The second competition requires her to make 3x + 2 cupcakes. Complete the expression to represent the total number of cupcakes. ______x + _______

OpenStudy (anonymous):

there you go:)

OpenStudy (anonymous):

I'm pretty sure it's 5x+6

OpenStudy (anonymous):

ok thanks anybody else have an input on it:)

OpenStudy (anonymous):

5x+6

OpenStudy (anonymous):

I belive you @aha.its.kristen

OpenStudy (anonymous):

ok thanks i have another

OpenStudy (anonymous):

Jon adds the expressions 5p + 2 and 2p + 1. Carlos adds the expressions 4p + 3 and 7p + 2. How much greater was Carlos’s sum than Jon’s sum? ________p + ________

OpenStudy (anonymous):

leave then come back in to see it better ok:)

OpenStudy (anonymous):

there you go:)

OpenStudy (anonymous):

7p+3 Jon / 11p+5 Carlous = 4p+2

OpenStudy (anonymous):

Carlos

OpenStudy (anonymous):

Jon's was 7p+3 and Carlos' was 11p+5 So subtract and you get 4p+2

OpenStudy (anonymous):

thanks 1 more:)

OpenStudy (anonymous):

hehe where so smart :) :p

OpenStudy (anonymous):

What is the perimeter of a triangle with the given side lengths? s1 = x – 3 ft s2 = x + 2 ft s3 = x + 4 ft P =______ x +_______ ft

OpenStudy (anonymous):

leave then come back inOk:)

OpenStudy (anonymous):

I think it's 3x+9

OpenStudy (anonymous):

@walter20

OpenStudy (anonymous):

3x+8 but i may be wrong

OpenStudy (anonymous):

:( what do i do:)

OpenStudy (anonymous):

im no good at perimeter

OpenStudy (anonymous):

Combining like terms There are 3 x's so it's 3x Then you add 3, 2, and 4. And you get 9

OpenStudy (anonymous):

ok:(

OpenStudy (anonymous):

add the side and figuer out x

OpenStudy (anonymous):

@aha.its.kristen What ddi you come up with???

OpenStudy (anonymous):

*did

OpenStudy (anonymous):

I passed thank you you guys:)

OpenStudy (anonymous):

80%

OpenStudy (anonymous):

4/5

OpenStudy (anonymous):

Was that right?^

OpenStudy (anonymous):

no it was 3x+3 it is ok though thank you soooooooooooooooooooooooooooooooooooooooooooooooooooooooo much:)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Yw

OpenStudy (anonymous):

@walter20 can you come to one of my rooms i am about to show:)

OpenStudy (anonymous):

ok